### v1

# PlagBench: Exploring the Duality of Large Language Models in Plagiarism Generation and Detection

Abstract
Recent literature has highlighted potential risks to academic integrity associated with large language models (LLMs), as they can memorize parts of training instances and reproduce them in the generated texts without proper attribution. In addition, given their capabilities in generating high-quality texts, plagiarists can exploit LLMs to generate realistic paraphrases or summaries indistinguishable from original work. In response to possible malicious use of LLMs in plagiarism, we introduce PlagBench, a comprehensive dataset consisting of 46.5K synthetic plagiarism cases generated using three instruction-tuned LLMs across three writing domains. The quality of PlagBench is ensured through fine-grained automatic evaluation for each type of plagiarism, complemented by human annotation. We then leverage our proposed dataset to evaluate the plagiarism detection performance of five modern LLMs and three specialized plagiarism checkers. Our findings reveal that GPT-3.5 tends to generates paraphrases and summaries of higher quality compared to Llama2 and GPT-4. Despite LLMs’ weak performance in summary plagiarism identification, they can surpass current commercial plagiarism detectors. Overall, our results highlight the potential of LLMs to serve as robust plagiarism detection tools.
PlagBench: Exploring the Duality of Large Language Models in Plagiarism Generation and Detection
Jooyoung Lee1      Toshini Agrawal2      Adaku Uchendu3 Thai Le4       Jinghui Chen1       Dongwon Lee1 1 The Pennsylvania State University, University Park, PA, USA, {jfl5838, jzc5917, dongwon}@psu.edu 2 Vellore Institute of Technology, Bhopal, India, agrawaltoshini@gmail.com 3 MIT Lincoln Laboratory, Lexington, MA, USA, adaku.uchendu@ll.mit.edu 4 Indiana University, Bloomington, IN, USA, tle@iu.edu
1 Introduction
Growing popularity of large language models (LLMs) has dramatically shifted the way users craft their writings. With one button click, LLMs such as OpenAI’s ChatGPT [^1] https://chatgpt.com and Google’s Gemini Team et al. (2023) can deliver high-quality texts per users’ requests. For instance, users can ask LLMs to generate a whole story or revise a draft writing. While an impressive ability, it is important to underscore that LLMs’ increasing applications in various writing tasks is accompanied by substantial risks to privacy, intellectual property (IP), academic integrity, and social ethics. Many recent papers (e.g., Carlini et al. (2021); Tirumala et al. (2022); Zhang et al. (2023); Carlini et al. (2023)) revealed that LLMs can memorize information from training data without overfitting, including sensitive information, and reproduce it during inference. Not limited to verbatim copies, they may emit paraphrased or elongated versions of training data Lee et al. (2023b). The observed behavior is problematic because: (1) LLMs are generally trained on vast amounts of text documents without considering how the content was originally released or whether it is copyrighted Brown et al. (2022); Henderson et al. (2023), and (2) LLMs are not currently equipped with the ability to accurately credit the original source of information as humans do McGowan et al. (2023); Huang and Chang (2023). Hence, if we compare these violations to human actions, they can be viewed as plagiarismthe act of employing another person’s work or ideas without proper attribution.
Figure 1: The overview of PlagBench construction processes and proposed RQs. Blue and red arrows denote the flow of RQ1 and RQ2, respectively.
To address plagiarism in LLMs, it is critical to improve automatic plagiarism detectors. Current detectors are typically trained on data related to human-authored plagiarism, while machine-generated plagiarism may exhibit significant differences Becker et al. (2023). Khalil and Er (2023) show that ChatGPT outputs were prone to bypass popular plagiarism check software Khalil and Er (2023). Automatic plagiarism detection task is also essential in a context where bad actors purposely utilize LLMs to obfuscate original content with paraphrases or summaries and evade existing detectors. For instance, Krishna et al. (2024); Sadasivan et al. (2023); Weber-Wulff et al. (2023) find that a simple paraphrase of machine-written text can fool widely used AI-generated content and plagiarism detectors. These results shed light on the urgent necessity of developing more reliable, robust, and universal plagiarism detectors.
To enable the development of modern LLM-aware plagiarism detectors, we posit that the curation of high-quality, large-scale, type-varying plagiarism datasets is necessary. Although there exist a few plagiarism detection corpora (e.g., Potthast et al. (2013); Wahle et al. (2021a)), they are not comprehensive and up-to-date to adequately reflect the rapidly evolving writing patterns influenced by LLMs. In this work, therefore, we first develop the PlagBench corpus, which consists of 46.5K text pairs covering three types of LLM-aware plagiarism (i.e., verbatim, paraphrase, and summary cases) in three writing domains (i.e., abstract of scholarly paper, story, news article). In the context of PlagBench, we investigate two research questions: (RQ1) how well can LLMs generate paraphrase and summary plagiarism of given texts? and (RQ2) how well can LLMs detect three types of plagiarism? Figure 1 illustrates PlagBench creation processes, as well as our research questions. Specifically, we use three state-of-the-art (SOTA) chat-based LLMs, Llama2-70b-chat Touvron et al. (2023), GPT-3.5 Turbo [^2] https://platform.openai.com/docs/models/gpt-3-5-turbo, and GPT-4 Turbo Achiam et al. (2023)) for PlagBench construction. Llama2-70b-chat, Llama3-70b-instruct, [^3] https://ai.meta.com/blog/meta-llama-3/ Mixtral-8x7B Jiang et al. (2024), GPT-3.5 Turbo, and GPT-4 Turbo are then used for the detection task. [^4] At the time of data generation, Llama3 models have not been released, and thus it was only employed in the detection process. Additionally, GPT-4o is excluded from both generation and detection tasks because it was unavailable during the time of the experiments.
Key findings of our experiments are as follows: (1) Among the three LLMs, GPT-3.5 Turbo produces the highest quality paraphrases and summaries, (2) LLMs like Llama3 and GPT-4, with just prompting, can outperform existing plagiarism checkers that are specifically trained for the task, and (3) LLMs generally have difficulty distinguishing summary plagiarism.
2 Related Work
2.1 Plagiarism Detection Corpus
The Microsoft Research Paraphrase Corpus (MRPC) Dolan and Brockett (2005) and PAWS Zhang et al. (2019b) are among the most renowned datasets for paraphrase identification. Yet, they are not suitable for plagiarism detection tasks because it only covers one type. Additionally, their samples are annotated at the sentence level, which limits their applicability in real-world scenarios where plagiarism in general occurs at the paragraph level. More recently, Wahle et al. (2021b) constructed a paragraph long machine-paraphrased plagiarism dataset using three neural language models such as BERT, RoBERTa, and Longformer. Following, the authors replaced these three models with more recently released LLMs, GPT-3, and T5, and produce machine-paraphrased texts Wahle et al. (2022). There are many independent machine-summarized text pairs (e.g., Cohan et al. (2018); Narayan et al. (2018)) as well, but there exists no comprehensive dataset covering both paraphrases and summaries generated by modern LLMs.
Potthast et al. (2013) introduced a novel plagiarism detection dataset with five obfuscation strategies: no obfuscation, random obfuscation, translation obfuscation, and summary obfuscation. These strategies utilized tools like Google Translate and text summarizers to mimic plagiarist behaviors. To the best of our knowledge, this is the first and only plagiarism detection dataset covering a wide range of plagiarism types. However, this dataset has not been updated since 2013.
2.2 Automatic Plagiarism Detection
There are two different formats of the automated plagiarism detection task: intrinsic and extrinsic plagiarism detection. While intrinsic plagiarism detection analyzes the document itself without resorting to the reference document, extrinsic plagiarism detection involves directly comparing the suspicious document against the reference to detect plagiarism instances Alzahrani et al. (2011). Intrinsic approaches use patterns such as linguistic differences within a document which indicates multiple authorship to detect plagiarism Gipp and Gipp (2014); Xiong et al. (2024); Potthast et al. (2010); Song et al. (2024); Stein et al. (2011). Due to high complexity of intrinsic plagiarism task, a majority of literature and software tools follow extrinsic plagiarism detection strategy. The extrinsic plagiarism detection task is generally divided into two subtasks: source retrieval (i.e., finding the most plausible pair of the source and the suspicious document) and plagiarism identification (i.e., classifying whether two documents are plagiarizing each other or not). Extrinsic approaches rely on similarity scores of two non-zero vectors transformed from the document pairs and apply certain thresholds Vani and Gupta (2017); Sağlam et al. (2024b); Potthast et al. (2010); Sağlam et al. (2024a); Lee et al. (2023a); Moravvej et al. (2023); Avetisyan et al. (2023). Our focus lies in the latter scope. We assume that we already have a text pair, the source document and the suspicious document obtained through PlagBench, and we concentrate on plagiarism identification.
3 Description of Automatic Plagiarism Detection
Suppose we have a source document $D_{src}$ and a suspicious document $D_{susp}$ that potentially plagiarizes $D_{src}$. Following the plagiarism detection task proposed by PAN, [^5] https://pan.webis.de we study three branches of plagiarism: verbatim, paraphrase, and summary. Verbatim plagiarism occurs when exact copies of words or phrases from $D_{src}$ are found in $D_{susp}$ without quotation marks. Paraphrase plagiarism, on the other hand, is when $D_{susp}$ rephrases $D_{src}$ using different words but retain the same meaning and structure without providing a citation. Synonymous substitution, word reordering and insertion/deletion are common paraphrasing techniques used by plagiarists Alvi et al. (2021). While both verbatim and paraphrase plagiarism cases tend to maintain the same length and the structure as the original text, summary plagiarism plagiarism involves succinctly summarizing $D_{src}$ to reuse its key points or ideas. As summaries leave out unnecessary information, they are prone to be significantly shorter than $D_{src}$. Unlike verbatim plagiarism, which can be effectively captured using simple string matching algorithms, paraphrase and summary plagiarism are more challenging to identify due to their limited lexical and syntactic similarities to the original text. Specifically, these two categories require careful attention to the meaning and context of the content rather than just high-level vocabulary overlaps.
4 PlagBench Creation
4.1 PlagBench Corpus Construction
Instruction tuning is a core training mechanism crucial for assisting LLMs in responding to human instructions. Once LLMs are pretrained with the objective of predicting subsequent words within large corpora, LLMs then undergo additional fine-tuning stage using diverse instruction datasets consisting of formatted instruction pairs. Instruction-tuned LLMs have demonstrated the ability to generalize to unseen instructions Raffel et al. (2020); Wei et al. (2021). Motivated by this, we leverage three prominent instruction-tuned LLMs, Llama2-70b-chat, GPT-3.5 Turbo, and GPT-4 Turbo, to create the PlagBench corpus. Our experiment design is as follows: given one source document $D_{src}$, three independent LLMs are prompted to rewrite it according to our description. We then consider their outputs as $D_{susp}$. Specifics of source document collection and suspicious document generation are described below.
4.1.1 Source Document Collection
Although plagiarism can occur in any type of writing, we target specific areas such as scholarly or creative works, where originality and intellectual property are highly valued. Li et al. (2023) present a collection of human-authored and machine-authored corpora from diverse writing tasks. Among their selections of corpora, we carefully choose three public English datasets: (1) SciXGen Chen et al. (2021) consisting of 200K+ abstracts of scientific articles (for scientific writing), (2) ROCStories Mostafazadeh et al. (2016) consisting of 50k five-sentence commonsense stories (for story writing); and (3) TLDR [^6] https://huggingface.co/datasets/JulesBelveze/tldr_news consisting of 7K+ TLDR tech newsletters (for news article writing). Their samples contain one human-written text and multiple machine-generated variations, all sharing either the same topic or text source as the human-written piece. For the purpose of the study, we only use the human-written text containing at least 5 sentences as the source text. [^7] According to our pilot experiment, paraphrases and summaries generated based on documents shorter than 5 sentences tend to too similar to distinguish even from human lens. We randomly select 2,400 samples from SciXGen and ROCStories corpora and 1,300 from TLDR for suspicious document generation. The sample size of TLDR is smaller than the other two datasets because content of TLDR tends to be short.
4.1.2 Suspicious Document Generation
Instruction-tuned LLMs have demonstrated strong abilities on text rephrasing Lingard (2023); Zhang et al. (2024); Tang et al. (2023). Hence, we employ Llama2-70b-chat, GPT-3.5 Turbo, and GPT-4 Turbo models to generate texts corresponding to two plagiarism labels (i.e., paraphrase and summary plagiarism) respectively. We do not use these models for verbatim plagiarism, as generating suspicious documents can be easily accomplished by simple copy-and-paste without any modification. Table 4 shows our hand-crafted prompt templates, each tailored to transform a single source document according to the provided task descriptions. As opposed to our expectations, our pilot experiment hints that some models, especially the Llama models, have difficulties in providing correct paraphrases or summaries on the fly. Thus, following findings that in-context learning is very effective in enabling LLMs to perform a new task without additional training Liu et al. (2021); Min et al. (2022), we provide three demonstrations, one from each writing domains, inside the prompt.
In plagiarism-free scenarios, existing datasets (e.g., Potthast et al. (2013), Foltỳnek et al. (2020)) paired random or dissimilar documents. However, this often results in pairs that are clearly different and thus easily identifiable. More challenging pairs are likely to occur when two documents share the same key topics and domain type, but their details do not overlap. Consequently, we instruct LLMs to write continuations based on provided keywords, domain information, and the first two sentences of the source document. Two meaningful keywords are automatically extracted using KeyBERT, [^8] https://maartengr.github.io/KeyBERT/index.html a package leveraging contextual embeddings from BERT Devlin et al. (2019) to generate keyphrases resembling the provided document. We do not perform in-context learning for plagiarism-free case generation because LLMs, originally trained for next token prediction, handled the task well without needing demonstrations.
We have in total 6,100 source documents across three writing domains from §4.1.1. Using three LLMs, we create three corresponding paraphrase, summary, and plagiarism-free documents per sample. For GPT-4, due to its relatively high API cost [^9] Llama2 families are free open-source models while GPT-3.5 turbo and GPT-4 Turbo charge \$3.00 and \$40.00 per 1M tokens respectively., we only utilize a subset (n=3,300) of source documents for generation. We use the version of gpt-3.5-turbo-1106 and gpt-4-1106-preview and apply temperature sampling (t = 0.9) for the generation configuration. As a result, the total number of generated suspicious documents is 46,500, a summation of 6,100 generations for SciXGen, 6,100 generations for ROCStories, and 3,300 generations for TLDR, multiplied by 3 plagiarism types.
4.2 PlagBench Quality Assurance
LLMs can occasionally fail to be faithful to the instruction and hallucinate Ji et al. (2023). For instance, when users instruct the LLM to paraphrase the given text, its output may contain sentences that are not properly paraphrased or are irrelevant to the original content. These hallucinated generations lead to an increase in false positives within PlagBench. To ensure the quality of our proposed dataset, we design detailed evaluation criteria and undergo rigorous filtering steps through both automatic and human evaluation. Automatic measures are intended to identify clearly incorrect cases, while human annotators can further refine the process by catching more complex errors that the automated methods might have overlooked. Table 2 describes 7 evaluation metrics used for paraphrase and summary plagiarism evaluation. These metrics are grounded on prior works studying evaluation methods for paraphrases Chen et al. (2020); Chen and Dolan (2011) or summaries Fabbri et al. (2021). For non-plagiarized cases, we use the same metrics; however, unlike in paraphrase and summary plagiarism, the suspicious documents should not satisfy any of them. We elaborate on how we assess the established metrics both automatically and manually in the following sections.
4.2.1 Assurance via Automatic Evaluation
As opposed to human evaluation, which is accurate but time-consuming, automatic measurement enables us to quickly and efficiently examine large volumes of data. Here we annotate all generated examples from paraphrase and summary plagiarism. Note that we do not perform automatic evaluation for non-plagiarism cases because it would overlap with our benchmark task and necessitate the use of the same plagiarism detectors we aim to evaluate. The evaluation process should take into account two pieces of text, one source document from §4.1.1 and one suspicious document from §4.1.2, and then assign numerical scores for corresponding aspects from Table 2 from one by one.
There has been continuous efforts in developing automated text quality estimation strategies. Among several existing metrics, we carefully choose a single representative metric for every aspect. For example, we use BERTScore Zhang et al. (2019a), one of the most common metrics for measuring semantic equivalence. It retrieves the token-level embedding using BERT and computes the summation of cosine similarities. It is shown to be more robust for paraphrase classification than other conventional metrics like BLEU Papineni et al. (2002) and METEOR Banerjee and Lavie (2005). For consistency measurement, we use AlignScore Zha et al. (2023) in which the authors train a information alignment model on 7 well-established tasks that include paraphrasing and summarization. The model returns a binary classification result (aligned v.s. not aligned). For language quality, we leverage the automated readability index Senter and Smith (1967) for fluency measurement and CoLAScore Zhu and Bhat (2020) for grammaticality, respectively.
Now we turn our attention to automatic summary evaluation. In the absence of gold-standard summaries for comparison with LLM-generated summaries, we rely on reference-free metrics. Relevance is then measured by BLANC Vasilyev et al. (2020), a novel approach that calculate the usefulness of a summary in helping BERT for language understanding task. This metric is ranked the highest in relevance alignment according to the Summeval benchmark Fabbri et al. (2021). For coherence measurement, we employ BARTScore Yuan et al. (2021), a comprehensive evaluation metric using a pre-trained sequence-to-sequence (seq2seq) model BART Lewis et al. (2020). We use the same models from paraphrase evaluation regarding consistency and language quality.
Most metrics, except for CoLAScore, provide continuous values, requiring threshold selection for filtering. The primary goal of automated evaluation is to eliminate obviously inaccurate pairs. Thus, we shuffle the order of suspicious document rows while keeping the source document order intact and create invalid pairs as a means of establishing a loose cut-off point. Refer to Appendix A.1 for experiment configurations and the threshold setup. We then exclude samples from a final set if the text pair does not satisfy all 3 or 4 aspects, depending on the plagiarism category. Lastly, to ensure that their generations do not overlap with verbatim plagiarism, we also compute the Levenshtein distance Miller et al. (2009) between two texts and remove near-duplicates. Among 31,000 (15,500 per plagiarism type) suspicious documents, 12,071 samples for paraphrase plagiarism and 13,445 samples for summary plagiarism remain intact. The document percentage breakdown per plagiarism type and domain type after automatic filtering is illustrated in Table 3.
Figure 2: Mean paraphrase evaluation aspect scores w.r.t. domain and model types (automatic evaluation)
Figure 3: Mean summary evaluation aspect scores w.r.t. domain and model types (automatic evaluation)
4.2.2 Assurance via Human Evaluation
Human annotation plays a critical role in verifying the accuracy and reliability of automated systems. We use Amazon Mechanical Turk (AMT) to hire human annotators for three separate annotation tasks: non-plagiarism evaluation, paraphrase evaluation, and summary evaluation. See Appendix A.2 for more details on the experiment design. Due to limited budgets, we perform manual annotation on 459 batches (i.e., 1,377 samples) for no plagiarism, 393 batches (i.e., 1,251 samples) for paraphrase plagiarism, and 417 batches (i.e., 1,251 samples) for summary plagiarism cases that are randomly sampled from §4.2.1. For data filtering, we follow the same evaluation criteria from our automated framework; if any of the evaluation aspects are flagged, we omit the particular samples. Among 3,879 documents, 1,358 samples from the non-plagiarism type, 1,014 from paraphrase, and 1,078 from summary plagiarism are deemed valid from a human perspective. In essence, nearly 89% of samples from automatic evaluation are considered reliable, suggesting a strong alignment between our automatic evaluation pipeline and human assessments.
Figure 4: Binary plagiarism detection (no plagiarism vs. plagiarism) performance of 5 LLMs w.r.t. prompt types. Dotted lines represent the performance of non-LLM based detectors.
5 Experimental Results and Discussion
5.1 RQ1. how well can LLMs generate paraphrase and summary plagiarism?
Since PlagBench is accompanied by both automatic and manual quality evaluation scores, we leverage them to conduct in-depth analyses on linguistic characteristics of paraphrases and summaries generated by three LLMs, Llama2-70b-chat, GPT-3.5 Turbo, and GPT-4 Turbo. To better understand natural phenomenon generated paraphrases and summaries before applying any filtering criteria, we solely rely on automated evaluation results. For paraphrases, we assess semantic equivalence, consistency, and language quality, while for summaries, we evaluate relevance, coherence, consistency, and language quality. Due to a large sample size, we randomly sample 1,000 subsets per writing domain in which each set contains on source document and three machine-rephrased versions and compare the mean of automatic evaluation scores.
Automatic Paraphrase Evaluation. Figure 2 shows mean evaluation scores with respect to types of models and writing domains in paraphrase generation. We exclude the gramaticality analyses since the majority of LLM-reworded texts are grammatically correct. models The results indicate that Llama2-70b-chat achieves significantly lower performance in BERTScore and Alignscore compared to both GPT-3.5 Turbo and GPT-4 Turbo. The observed pattern persists across three writing domains. Particularly, Llama2-70b-chat suffers from generating factually consistent paraphrases for short stories (ROCStories). This finding is consistent with existing literature Mishra et al. (2024); Dahl et al. (2024) highlighting the vulnerability of Llama2 in hallucination. Within the scope of semantic equivalence and consistency metrics, GPT-3.5 Turbo achieves the highest score, even occasionally better than GPT-4 Turbo. Analyses on readability score [^10] We use the automated readability index to measure readability. A higher readability index hints that the text is more complex and may necessitate a higher level of education or reading proficiency to comprehend. gaps between the human-written source text and the machine-generated suspicious text suggest that all three models significantly complicate ROCStories during rephrasing. In regards to SciXGen and TLDR, GPT models tend to produce more sophisticated texts, whereas Llama2-70b-chat often degrades linguistic quality. Substantial readability gaps, especially in the GPT-4 version, are somewhat anticipated, as Onder et al. (2024); Momenaei et al. (2023) have noted that ChatGPT’s outputs can be challenging to understand, often demanding a college-level proficiency in linguistic skills.
Automatic Summary Evaluation. Summary evaluation results are illustrated in Figure 3. Regardless of writing domain categories, GPT-3.5 Turbo is ranked the lowest in BARTScore, followed by Llama2-70b-chat and and then GPT-4 Turbo. The lower the BARTScore the more coherent given a pair of the source document and the machine-summarized document. There is no noticeable differences between three models regarding the Alignscore, suggesting that they excel at producing factually consistent summaries. Still, GPT3-3.5 Turbo outperforms both models in the domains for SciXGen and TLDR. Also, BLANC scores are shown to be the highest for GPT-4 Turbo while GPT-4 Turbo was ranked the lowest. The higher BLANC score is the more relevant and useful the given summary is. As opposed to paraphrase analyses which take into account the difference between two pieces of text, here we only compute the readability index scores of LLM-generated summaries. This is because we do not have reference summaries to compare against. Consistent with paraphrase generation, GPT-4 Turbo tends to produce summaries with higher readability scores in comparison to Llama2-70b-chat and GPT-3.5 Turbo. Interestingly, summaries generated by Llama2-70b-chat are associated with higher readability scores than GPT-3.5 Turbo. Based on thes findings, we conclude that GPT-3.5 Turbo is generally the best paraphraser and summarizer in terms of all aspects, except for fluency/readability. Also, while Llama2-70b-chat is ranked the worst model for paraphrasing task, GPT-4 Turbo suffers the most for providing high-quality summaries.
N/A Llama2-70b-chat GPT-3.5 Turbo GPT-4 Turbo
Detectors Verb No Para Summ No Para Summ No Para Summ Acc
Llama2-70b-chat
66.67% 36.29% 68.88% 6.66% 30.37% 55.55% 8.88% 33.33% 71.11% 0% 39.5%
Llama3-70b-instruct
100% 85.92% 95.55% 11.11% 81.48% 100% 11.11% 78.51% 100% 8.88% 75.8%
Mixtral 8x7B-instruct
88.14% 34.81% 55.55% 11.11% 29.62% 66.66% 8.88% 21.48% 77.77% 11.11% 41.85%
GPT-3.5 Turbo
77.77% 47.41% 86.66% 17.77% 42.22% 97.77% 2.22% 41.48% 91.11% 2.22% 51.35%
GPT-4 Turbo
99.25% 93.33% 82.22% 17.77% 92.59% 100% 15.55% 85.18% 100% 15.55% 80.12%
GPTZero 37.03% 100% 6.66% 0% 100% 2.22% 4.44% 100% 0% 0% 56.91%
Prepostseo 5l.85% 91.85% 20% 0% 96.29% 17.77% 2.22% 93.33% 4.44% 6.66% 58.39%
WWW’23 detector 82.22% 80% 88.88% 82.22% 77.77% 86.66% 84.44% 83.7% 77.77% 91.11% 82.34%
Table 1: Plagiarism type classification (original (i.e., ‘No’) vs. verbatim (i.e., ‘Verb’) vs. paraphrase (i.e., ‘Para’) vs. summary (i.e., ‘Summ’) plagiarism) performance of 8 detectors w.r.t. the models used for generation. For non-LLM approaches, we compute the category-wise breakdown from binary classification as they are not suitable for this task. ‘Acc’ in the leftmost column indicates the overall accuracy regardless of category types. The highest values among LLM-based approaches are highlighted in red, while the highest values for non-LLM-based approaches are highlighted in blue.
5.2 RQ2. how well can LLMs detect plagiarism?
Experimental Setup. To investigate LLMs’ capabilities in plagiarism detection, we present two tasks: 1) binary classification (plagiarism-free vs. plagiarism) and 2) plagiarism type identification (plagiarism-free vs. verbatim vs. paraphrase vs. summary plagiarism). Unlike RQ1, we leverage human annotation results to curate the evaluation data. Specifically, we randomly sample 810 pairs of $D_{susp}$ and $D_{susp}$ from human annotation subset of PlagBench. The breakdown of the 810 text pairs consists of 45 text pairs from no plagiarism labels and 45 from three plagiarism labels, generated by 3 models within 3 different domains. We experiment with four prompt variations: zero-shot vanilla prompting, zero-shot Chain-of-Thought (CoT) prompting Wei et al. (2022), few-shot vanilla prompting, and few-shot CoT prompting. CoT prompting is one of the most popular prompt techniques to enhance models’ downstream task performance by eliciting reasoning. As shown in Table 5, we incorporate our defined plagiarism categories and their respective definitions into the prompt, aiming to enhance the alignment of LLMs with our task. For few-shot experiments, we provide six demonstrations consisting of a mix of three plagiarism pairs and three plagiarism-free pairs, ensuring no overlap with the evaluation data. We leverage GPT-4 to obtain high-quality reasoning for few-shot CoT prompting.
Five LLMs are tested in total: Llama2-70b-chat, Llama3-70b-instruct, Mixtral-8x7B-instruct, GPT-3.5 Turbo, and GPT-4 Turbo. We employ greedy decoding (i.e., choosing the most plausible token at each generation step) for answer prediction because it is widely used for various detection tasks and supports the reproducibility of the results to some extent.
We additionally incorporate three non-LLM-based plagiarism detectors as baseline models. Two of these, GPTZero [^11] https://gptzero.me/ and Prepostseo, [^12] https://www.prepostseo.com/plagiarism-checker are commercial plagiarism checkers, while the third is a detector proposed by Lee et al. (2023b). We denote Lee et al. (2023b)’s detector as the WWW’23 detector for the rest of the paper. All these tools rely on semantic similarity measures and text alignment algorithms to distinguish plagiarism.
Binary Plagiarism Detection Results. Figure 4 shows the plagiarism detection performance of $D_{src}$ and $D_{susp}$ in a binary setting (original vs. plagiarism). We report accuracy scores since the label distribution is balanced; 405 examples out of 810 are plagiarism-free cases, and the remaining 405 examples belong to one of three plagiarism types. The results of zero-shot vanilla prompting, the most basic setup, hint that Llama2-70b-chat and Mixtral-8x7B-instruct perform poorly with accuracy scores worse than a random guess (Acc = 0.5). According to our qualitative inspection, the errors from Llama2-70b-chat and Mixtral-8x7B-instruct arise because their final outputs do not include a binary prediction. 26.7% of Llama2-70b-instruct’s output and 11.6% of Mixtral-8x7B-instruct’s output do not provide the prediction at all. Additionally, their predictions tend to be highly skewed to the negative class. In contrast, the performance of Llama3-70b-instruct is twice higher than that of Llama2-70b-chat. Similarly, there is a significant performance improvement from GPT-3.5 Turbo to GPT-4 Turbo. These findings are reasonable in a sense that Llama3 and GPT-4 are upgraded versions of Llama2 and GPT-3.5, respectively. A similar phenomenon has been reported in the task of medical final examination Rosoł et al. (2023). CoT prompting and few-shot learning tends to significantly boost detection performance across all models, consistent with previous literature Wei et al. (2022); Brown et al. (2020). Especially, Mixtral-8x7B-instruct and GPT-3.5 Turbo demonstrates a noticeable performance gain from few-shot CoT prompting. We inspect that demonstrations of different plagiarism categories and the corresponding rationale that supports the correct answer assist them to identify more complicated plagiarism categories like paraphrases and summaries. The performance gain of Llama3-70b-instruct and GPT-4 Turbo obtained from few-shot CoT prompting is not as significant as that of other models, but still these achieve near-perfect performance. Commercial plagiarism detectors, GPTZero and Prepostseo, is shown to achieve roughly 10% higher performance than random guessing. They outperform three LLMs in a zero-shot setting, but their performance is substantially lower than recently published LLMs such as Llama3-70b-instruct and GPT-4 Turbo. This may be due to these tools relying on simple string-matching algorithms designed specifically for detecting verbatim plagiarism, rather than accommodating a broader range of plagiarism categories. Conversely, the WWW’23 detector exhibits stronger performance as it is specifically tailored for this task. Nonetheless, four LLMs have been shown to surpass its performance through few-shot CoT prompting. Overall, these results highlight the potential of LLMs in effectively detecting plagiarism.
Plagiarism Type Classification Results. Table 1 shows detection results of 8 detectors on plagiarism type classification depending on plagiarism types and models used for generation. Of the four prompting techniques, we resort to zero-shot CoT prompting for this experiment. This decision is motivated by the fact that its inference time is significantly faster than few-shot CoT prompting with improved detection performance. Equivalent to binary plagiarism detection performance measurement, we report accuracy scores. In line with the results from binary Plagiarism Detection, Llama3-70b-instruct and GPT-4 Turbo are the highest performing LLMs. In particular, their identification of verbatim and paraphrase plagiarism achieve 99-100% accuracy. GPT-3.5 Turbo is demonstrated to be quite capable of distinguishing paraphrase plagiarism, but it performs poorly in identifying plagiarism-free content and summarized content. Most notably, LLMs perform poorly in detecting summary plagiarism across all genres. Most of their errors stem from short texts being confused for paraphrase plagiarism instead of summary plagiarism. Now, looking at traditional plagiarism checkers, we find that GPTZero and Prepostseo excel at identifying non-plagiarized content compared to the WWW’23 detector. Yet, the WWW’23 detector surpasses them in paraphrase and summary detection. This discrepancy may be due to differing definitions of plagiarism; since these tools are designed for detecting verbatim plagiarism, they may incorrectly classify text pairs with minimal lexical overlap as non-plagiarized.
6 Conclusion
We present a novel large-scale plagiarism detection benchmark, PlagBench: a collection of 46.5K artificial plagiarism cases generated by three cutting-edge LLMs. The development and analysis of the PlagBench dataset underscores the nuanced challenges and opportunities presented by LLMs in combating plagiarism while also acknowledging their potential vulnerabilities to being used for unethical practices. We envision PlagBench as a universal standard for evaluating plagiarism detection methods.
Limitations
The current research has two limitations. First, we use a single manually crafted prompt template for generating machine-plagiarized documents and detecting plagiarism. There are, however, various prompting techniques and strategies for automatically optimizing prompts. We expect future work to explore a broader range of prompts for both generation and detection purposes. Second, our experiments are based on instruction-tuned decoder-only LLMs. It is uncertain whether the observed findings would apply to other types of LLMs with different architectures. Future research should revisit our RQs using more diverse model types.
Ethics Statement
This research involves the use of LLMs to simulate plagiaristic behavior. All source documents utilized in this study are derived from publicly accessible open-source datasets. To promote transparency and reproducibility, we will make all data and code used in our experiments available to the research community. We acknowledge the potential ethical concerns associated with generating synthetic plagiarism documents. Therefore, we emphasize that these documents should be used exclusively for research purposes aimed at understanding and mitigating plagiarism. We advise against any misuse of the synthetic data, such as using it to deceive or harm others. Our work is intended to contribute to the development of more effective plagiarism detection tools and to enhance academic integrity. By sharing our resources, we aim to support the broader research community in these endeavors.
Regarding human annotation, our research protocol was approved by the Institutional Review Board (IRB) at our institution. We only recruited annotators that are 18 years old or over. All annotators were paid over minimum wage rate.
Appendix A Appendix
A.1 Automatic Evaluation of LLM-Generated Plagiarism
We leverage publicly available two datasets, PAWS Zhang et al. (2019b) and SummEval Fabbri et al. (2021), to find the optimal thresholds for data filtering regarding paraphrase evaluation and summary evaluation, respectively. Both datasets contain human annotation labels signaling whether a pair of texts is correct paraphrases or not. PAWS provide one binary label for each text pair, whereas SummEval contains a list of 8 annotation results in which the scores range from 1 to 5 evaluated across 4 dimensions: coherence, consistency, fluency, and relevance.
The threshold selection process is straightforward; we apply our established automated evaluation measurements to the subset of these two corpora and compute their mean scores of bad examples. Specifically, we compute mean scores of entries with ‘0’ label from the PAWS dataset and entries with average annotation scores lower than 2.5 from the SummEval dataset. These mean scores are then used as a reference to assign filtering thresholds. Give two pieces of evaluation texts $(d,d^{*})$, let’s denote paraphrase evaluation results as $R_{\text{para\_eval}}$. $R_{\text{para\_eval}}$ can be expressed as below:
$R_{\text{para\_eval}}=\begin{cases}1&\text{if }\text{Alignscore}(d,d^{*})>0.5\\ &\quad\land\text{BERTScore}(d,d^{*})>0.5\\ &\quad\land\text{Readability\_Gap}(d,d^{*})>-0.5\\ &\quad\land\text{COLAscore\_Improv}(d,d^{*})\geq 0\\ 0&\text{else}.\end{cases}$
Summary evaluation results are denoted as $R_{\text{summ\_eval}}$.
$R_{\text{summ\_eval}}=\begin{cases}1&\text{if }\text{Alignscore}(d,d^{*})>0.5\\ &\quad\land\text{BLANC}(d,d^{*})>0.0\\ &\quad\land\text{BARTScore}(d,d^{*})>-4.0\\ &\quad\land\text{Readability}(d,d^{*})>10\\ &\quad\land\text{COLAscore\_Improv}(d,d^{*})\geq 0\\ 0&\text{else}.\end{cases}$
Given these two definitions, we omit all samples associated with ‘0’ in regards to $R_{\text{para\_eval}}$ and $R_{\text{summ\_eval}}$.
A.2 Details on Human Evaluation of LLM-Generated Plagiarism
Each task consists of 3 batches of one source text and three suspicious texts generated by Llama2-70b-chat, GPT-3.5 Turbo, and GPT-4 Turbo. Given a pair of the source text and machine-paraphrased or machine-summarized texts, we first verbally describe our established evaluation aspects on the task description and ask them to answer if the described aspect is satisfied or not. For plagiarism-free cases, the definitions of paraphrase and summary plagiarism were provided, and annotators are tasked to identify two texts belong to either paraphrase or summary plagiarism. See Figure 6, 7, and 5 for the user study interface design. To ensure the quality of responses submitted by annotators, we enforce four built-in worker qualifications, involving(1) HIT Approval Rate of ${\leq}98\%$, (2) a minimum of 1,000 Approved HITs, (3) Masters qualification status, and (4) U.S-based workers. Moreover, we purposely add one dummy question per batch and reject the response providing incorrect answers to the question. Given that the estimated completion time per batch is 10 minutes, workers are compensated \$$\$1.5$ per batch based on United States average hourly wages. We hire 3 annotators per sample in order to assign majority voting results as gold labels.
A.3 Model Size And Budget
We enlist model size used for our experiment here: for Llama2 and Llama3, we use the largest scale among their models, which is 70B. GPT-3.5 is expected to contain 175B parameters, equivalent to GPT-3. Meanwhile, GPT-4 has 1.76T parameters, and Mixtral has 46.7B total parameters. For all experiments, we leverage API calls from Huggingface and OpenAI official website.
Figure 5: AMT experiment interface example for plagiarism-free case evaluation
Figure 6: AMT experiment interface example for paraphrase plagiarism evaluation
Figure 7: AMT experiment interface example for paraphrase plagiarism evaluation
Type Metric Description
Paraphrase
Semantic
equivalence
Evaluates whether the paraphrased text conveys the same meaning as the source text.
Consistency
Evaluates whether the paraphrased text accurately represents the information contained
in the source text without the inclusion of errors or distortions.
Language
quality
Evaluates whether the paraphrased text maintains or improves upon the quality of the
source text in regards to fluency and grammaticality.
Summary Relevance
Evaluates whether the summary covers all the essential information and key points from
the source text while omitting irrelevant or tangential details.
Coherence
Evaluate whether the summary is well-structured and not just a random assortment of
information from the source text.
Consistency
Evaluates whether the summary accurately represents the content and meaning of the
source text without the inclusion of errors or distortions.
Language
quality
Evaluates whether the summary maintains or improves upon the quality of the source
text in regards to fluency and grammaticality.
Table 2: Descriptions of evaluation metrics for paraphrase and summary Plagiarism.
Domain Model # of document pairs
# of remaining pairs
# of remaining pairs
SciXGen Llama2-70b-chat 2,400 1,179 (49.12%) 2,147 (89.46%)
GPT-3.5 Turbo 2,400 1,764 (73.5%) 2,379 (99.12%)
GPT-4 Turbo 1,300 1,085 (83.46%) 1,266 (97.38%)
ROCStories Llama2-70b-chat 2,400 1,699 (70.79%) 1,975 (82.29%)
GPT-3.5 Turbo 2,400 2,385 (99.38%) 1,614 (67.25%)
GPT-4 Turbo 1,300 1,295 (99.62%) 1,166 (89.69%)
TLDR Llama2-70b-chat 1300 753 (57.92%) 1,060 (81.54%)
GPT-3.5 Turbo 1300 1,230 (99.38%) 1,181 (90.85%)
GPT-4 Turbo 700 681 (97.29%) 657 (93.86%)
Table 3: Dataset statistics after automatic filtering for paraphrase and summary plagiarism. The percentage in the bracket represents the percentage of remaining document pairs remaining out of original pairs.
Plagiarism category Template
Paraphrase
Paraphrase the following text while keeping its meaning.
Text: {source_doc_demonstration1}
Paraphrased: {paraphrase_demonstration1}
Text: {source_doc_demonstration2}
Paraphrased: {paraphrase_demonstration2}
Text: {source_doc_demonstration3}
Paraphrased: {paraphrase_demonstration3}
Text: {target_source_doc}
Paraphrased:
Summary
Summarize the following text in 1-3 sentences.
Text: {source_doc_demonstration1}
Summarized: {summary_demonstration1}
Text: {source_doc_demonstration2}
Summarized : {summary_demonstration2}
Text: {source_doc_demonstration3}
Summarized: {summary_demonstration3}
Text: {target_source_doc}
Summarized:
Plagiarism-Free
Based on the provided text passage and keywords, write its continuation in a style of {genre}.
When generating, make sure that the continuation is coherent while including all keywords.
Text: {target_sentences}
Keywords: {keywords}
Continuation:
Table 4: Prompt templates for plagiarism generation task. Text in violet represents instructions and text in gray represents few-shot demonstrations. Lastly, text in magenta is used for the actual task.
Prompt type Template
Zero-shot vanilla
prompting
(binary detection)
There are three types of plagiarism: • Verbatim plagiarism: the evaluation text consists of exact copies of words or phrases without transformation from the source text without citation. • Paraphrase plagiarism: the evaluation text is rephrased or rewritten using different words but retain the same meaning and structure as the source text without citation. • Summary plagiarism: the evaluation text encapsulates the most essential points of the source text into a shorter form without citation. If the evaluation text does not belong to any of three plagiarism categories, it means plagiarism-free. Given a pair of text and provided plagiarism definitions, does the evaluation text plagiarize the source text (yes/no)? Please format your final response as ‘Answer: {{response}}’. Source text: {source_doc} Evaluation text: {susp_doc}
Zero-shot CoT
prompting
(binary detection)
There are three types of plagiarism:
• Verbatim plagiarism: the evaluation text consists of exact copies of words or phrases
without transformation from the source text without citation.
• Paraphrase plagiarism: the evaluation text is rephrased or rewritten using different words but
retain the same meaning and structure as the source text without citation.
• Summary plagiarism: the evaluation text encapsulates the most essential points of the source text
into a shorter form without citation.
If the evaluation text does not belong to any of three plagiarism categories, it means plagiarism-free.
Given a pair of text and provided plagiarism definitions, does the evaluation text plagiarize the source text
(yes/no)? First think step-by-step and format your final response as ‘Final answer: {{response}}’.
Source text: {source_doc}
Evaluation text: {susp_doc}
Answer: Let’s think step-by-step.
Zero-shot vanilla
prompting
(type identification)
There are three types of plagiarism:
• Verbatim plagiarism: the evaluation text consists of exact copies of words or phrases
without transformation from the source text without citation.
• Paraphrase plagiarism: the evaluation text is rephrased or rewritten using different words but
retain the same meaning and structure as the source text without citation.
• Summary plagiarism: the evaluation text encapsulates the most essential points of the source text
into a shorter form without citation.
• No plagiarism: the evaluation text does not belong to any of three plagiarism categories.
Given a pair of text and provided plagiarism definitions, what type of plagiarism
(no/verbatim/paraphrase/summary) does the evaluation belong to when compared to the source text?
Please format your final response as ‘Answer: {{response}}’.
Source text: {source_doc}
Evaluation text: {susp_doc}
Zero-shot CoT
prompting
(type identification)
There are three types of plagiarism:
• Verbatim plagiarism: the evaluation text consists of exact copies of words or phrases
without transformation from the source text without citation.
• Paraphrase plagiarism: the evaluation text is rephrased or rewritten using different words but
retain the same meaning and structure as the source text without citation.
• Summary plagiarism: the evaluation text encapsulates the most essential points of the source text
into a shorter form without citation.
• No plagiarism: the evaluation text does not belong to any of three plagiarism categories.
Given a pair of text and provided plagiarism definitions, what type of plagiarism
(no/verbatim/paraphrase/summary) does the evaluation belong to when compared to the source text?
First think step-by-step and
format your final response as ‘Answer: {{response}}’.
Source text: {source_doc}
Evaluation text: {susp_doc}
Answer: Let’s think step-by-step.
Table 5: Prompt templates for plagiarism detection task.